OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify sqrt320 and provide the simplified form.

OpenStudy (anonymous):

\[8\sqrt{?}\] is my answer, i know that :)

OpenStudy (anonymous):

\[8\sqrt{5}\]

OpenStudy (anonymous):

take the terms that repeat themselves twice outside the square root sign. But before that u will have to write the factors of the number 320 320=2*2*2*2*2*2*5